What are the foci of the ellipse 17x2 + 8y2 = 136? show work please
Answer by lwsshak3(11628)

What are the foci of the ellipse 17x2 + 8y2 = 136? show work please
**
17x^2+8y^2=136
divide by 136
x^2/8+y^2/17=1
This is an ellipse with vertical major axis of the standard form: (x-h)^2/b^2+(y-k)^2/a^2=1, a>b
for given equation:
Center:(0,0)
a^2=17
b^2=8
c^2=a^2-b^2=17-8=9
c=√9=3
Foci: (0,0±c)=(0,0±3)=(0,3) and (0,-3)